OpenStudy (anonymous):

If a^2 = by + cz, b^2 = cz +ax, c^2 = ax +by then the value of x/(a+x) + y/(b+y) + z/(c+z) = a.) a+b+c, b.) (1/a) + (1/b) + (1/c) c.) 1 d.) 0

OpenStudy (anonymous):

now multiply both num and denom of x/(a+x) by a we then have ax/(a^2 +ax) we know a^2 = by + cz hence ax/(a^2 +ax) becomes ax/( by + cz +ax)

OpenStudy (anonymous):

cna you do it further

OpenStudy (anonymous):

now multiply both num and denom of y/(b+y) by b and that of z/(c+c) by c ..

OpenStudy (anonymous):

x/(a+x) + y/(b+y) + z/(c+z) =ax/(a^2 +ax) + by/(b^2+by) + cz/(c^2+cz)

OpenStudy (anonymous):

use a^2 = by + cz, b^2 = cz +ax, c^2 = ax +by then we have ax/(ax+by+cz) + by/(ax+by+cz) +cz/(ax+by+cz) =(ax+by+cz)/(ax+by+cz) =1

OpenStudy (anonymous):

Thank you! Really appreciated...